# Stochastic and incremental subgradient methods for convex optimization on Hadamard spaces

### v3

@@ 1 Introduction
The algorithm then tracks the geodesic ray from $x$ associated with $b$ at speed $s$ for a judiciously chosen time interval, before repeating the process. Analogous subgradient inequalities appeared in [26, 28], and indeed, the notion of a Busemann subgradient that we introduce is related to ideas in [20, 28]. The algorithms we present generalize the classical stochastic subgradient method, when choosing components uniformly at random, and the incremental subgradient algorithm of [30]: furthermore, they enjoy the same complexity. While the $p$-mean problem serves as a familiar testing ground for our algorithms, we emphasize that Busemann subgradient ideas apply more broadly: examples include the minimum enclosing ball problem [3], and, as highlighted recently in [16] [^1] While revising this paper from the original posted manuscript [21], we learned of subsequent developments involving the Busemann subgradient idea [16]., Tyler’s M-estimator [38] and Horn’s problem [27].
@@ 2 Geodesic geometry and convexity
We will restrict our attention to a class of functions that interacts nicely with the geometry of the given Hadamard space in a way that goes beyond geodesic convexity. Fundamental to our development is an appropriate notion of direction in a Hadamard space; in the Euclidean setting, the role played by directions and the associated compactification of $\mathbb{R}^{n}$ has been emphasized as a pillar of modern variational analysis [35, Chapter 3].
@@
We first review some standard tools to study the geometry at infinity of a Hadamard space, a more detailed discussion of which can be found in [13, Chapter II.8]. Two rays $r,r^{\prime}$ are said to be asymptotic if there exists a positive constant $K\geq 0$ such that $d(r(t),r^{\prime}(t))\leq K$ for all $t\geq 0$. This defines an equivalence relation on rays in $X$: the set of equivalence classes is denoted by $X^{\infty}$ and called the boundary of $X$ at infinity, or simply the boundary. Note our assumption that $X$ has at least two points and the geodesic extension property guarantees the boundary is nonempty. The equivalence class of a particular ray $r$ is denoted by $r(\infty)$, and we say a ray $r$ has direction $\xi\in X^{\infty}$ if $r$ belongs to the equivalence class $\xi$. Given $x\in X$ and $\xi\in X^{\infty}$, there exists a unique ray $r$ issuing from $x$ such that $r(\infty)=\xi$ [13, II.8.2]. If we fix an arbitrary reference point $\bar{x}\in X$, we can thus identify any $\xi\in X^{\infty}$ with the unique ray $r_{\bar{x},\xi}$ issuing from $\bar{x}$ with direction $\xi$, and hence to the unique corresponding Busemann function which we may denote by either $b_{r_{\bar{x},\xi}}$ or $b_{\bar{x},\xi}$. To avoid cumbersome notation, we henceforth fix such a reference point $\bar{x}\in X$ and denote the corresponding Busemann function $b_{r_{\bar{x},\xi}}$ by simply $b_{\xi}$. We note the following fact [13, Corollary II.8.20].
@@
The space $X^{\infty}$ is naturally endowed with the so-called cone topology [13, Definition II.8.6], and notably this space is first-countable [13, Proof of Theorem II.8.13]. In particular, the cone topology is completely specified by the convergent sequences, so let us mention that $\xi_{n}\to\xi\in X^{\infty}$ if and only if $b_{\xi_{n}}\to b_{\xi}$ uniformly on bounded subsets of $X$. The latter property, and hence the cone topology, is independent of the implicit basepoint (cf. [13, Proposition II.8.8]), due to Proposition 2.1.
If $X$ is proper then $X^{\infty}$ is compact [13, Definition II.8.6]. Our interest in topologizing $X^{\infty}$ comes from optimization: it will be desirable to have a large class of compact subsets of $X^{\infty}$ so that certain continuous functions defined on $X^{\infty}$ will attain their maximum. In the literature, the set $X^{\infty}$ is often endowed with a stronger metric topology induced by an angular metric (see for example [13, Proposition II.9.7], [23]). This angular metric is convenient for the study of large-scale geometry of the space $X$ but is often too strong to be useful for our purposes. For example, the boundary at infinity of the $n$-dimensional hyperbolic space $\mathbb{H}^{n}$ is discrete in this angular metric, whereas it is homeomorphic to $\mathbb{S}^{n-1}$ in the cone topology (see Example 2.3). In particular, the only compact sets are finite when the angular metric is used. Henceforth we will only make use of the cone topology.
Convergence in $X^{\infty}$ can be understood more geometrically via pointwise convergence of rays, a fact stated in [17, p.7] without proof, and also noted in [4, p.29]. The next proposition is a precise formulation: for completeness, we include a proof in the appendix. The proof and result resemble [13, Proposition II.8.19], but that result concerns unbounded sequences in $X$ that converge to points in $X^{\infty}$ whereas we emphasize sequences converging in $X^{\infty}$ itself.
@@ 3 Busemann subgradients and envelopes
is the Busemann function for the ray $r$. The argument on the right-hand side is nonincreasing in $t$, by [13, Lemma II.8.18(1)], and equals d(y,a)d(x,a)d(y,a)-d(x,a) when $t=d(x,a)$, so the result follows. ∎
@@
We consider $X=\mathbb{H}^{2}$, The subset $F=\left\{(x,y)\in\mathbb{H}^{2}\mid y=0\right\}$ is closed and geodesically convex (it is a geodesic line in $\mathbb{H}^{2}$), so the function $f(p)=\operatorname{dist}(p,F)$ is geodesically convex ([13, II.2 Corollary 2.5]). However, $f$ is not Busemann subdifferentiable on any neighborhood of $F$. To see this, let $U$ be a neighborhood of $F$, consider any $(x_{0},y_{0})\in U$ with $y_{0}\neq 0$, and suppose there exists a Busemann subgradient $\xi,s]$ at $(x_{0},y_{0})$, with $s>0$ because $(x_{0},y_{0})$ does not minimize $f$. We can identify $\xi\in X^{\infty}$ with a point $\xi=(\xi_{1},\xi_{2})\in\mathbb{S}^{1}$, and we will use this to write the Busemann function $b_{\xi}$ explicitly. By definition $f-\left\langle\cdot,[\xi,s]\right\rangle$ is minimized at $(x_{0},y_{0})$, i.e.
@@
Choosing our reference point to be $(\bar{x},\bar{y})=(0,0)$, [13, II.8.24(2)] tells us the Busemann function $b_{\xi}$ has the following form:
@@
(Continuous-compact-representable functions) Endowing the boundary $X^{\infty}$ with the cone topology opens up the following natural construction. Suppose $K\subseteq CX^{\infty}$ is compact and $g\colon CX^{\infty}\to(-\infty,+\infty]$ is continuous on $K$ and $\infty$ outside of $K$. Because the pairing is also continuous (Proposition 2.3(1)), the function $g^{\circ}$ is a Busemann envelope, and hence Busemann subdifferentiable by Proposition 3.4. This construction generates a rich class of examples. In particular, in Euclidean space we recover the class of “almost sublinear” functions: those functions $f$ such that $\sup_{x\in\mathbb{R}^{n}}|f(x)-h(x)|<\infty$ for some sublinear function $h$ [11, Proposition 4.5]. This class covers all convex optimization problems with compact feasible regions, by extending the objectives to have this growth property outside the feasible region. More generally, consider the case when the space $X$ is proper, and hence the boundary $X^{\infty}$ is compact. Then compact subsets of the boundary cone $CX^{\infty}$ are abundant, because if $F\subseteq X^{\infty}$ is closed and $I\subseteq\mathbb{R}_{+}$ is compact, then the image of $F\times I$ under the quotient map is compact.
@@
which holds with equality if and only if $\xi,s]$ is a Busemann subgradient for $f$ at $x$, analogously to the classical case. The only difference between the definitions of $f^{*}$ and $f^{\bullet}$ is the sign of the pairing term. For comparison, [8, Definition 1] concerns a notion of Fenchel conjugate on Hadamard manifolds with the sign on the pairing term matching inequality (14).
@@ 4 The geometry of Busemann subdifferentiability
Note $s_{k}>0$ because $x_{k}$ does not minimize $f$, hence $b_{x_{k},\xi_{k}}(y)\leq 0$ for all $y\in F$. Properness of $X$ implies that $X^{\infty}$ is compact, allowing us to extract a convergent subsequence $\xi_{k}\to\xi$. The map $(x,\xi)\mapsto b_{x,\xi}(y)$ is continuous [4, II.1] so passing to the limit implies $b_{x,\xi}(y)\leq 0$ for $y\in F$. Thus $F$ admits a supporting horoball at each boundary point as desired. ∎
@@
Just as in Euclidean space, continuous convex functions on the space $X$ must be locally Lipschitz, by [28, Lemma 4.3]. The following result relates Busemann subgradients to the local Lipschitz constant: for an analogous Euclidean result, see [7, Theorem 3.61].
@@
(ii) Suppose $f$ is $L$-Lipschitz on $C$ and take any Busemann subgradient $\xi,s]$ at $x\in C$. If $s=0$ there is nothing to prove, so we assume $s>0$. Let $r$ be the ray issuing from $x$ with direction $\xi$. Since $X$ has the geodesic extension property $r$ can be extended to a geodesic line $\tilde{r}\colon\mathbb{R}\to X$ by [13, Lemma II.5.8(2)]. The geodesic line $\tilde{r}$ is continuous with $\tilde{r}(0)=x\in C$, and $C$ is open so there exists $\varepsilon>0$ such that $\tilde{r}(-\varepsilon)\in C$. By (6) we have
@@
This space arises as a subspace of the tree space $\mathcal{T}_{4}$ (see Section 8), and is illustrated in [10, Figure 12] as well as Figure 4.1. We will revisit this space when discussing the Busemann subdifferentiability of a sum of Busemann subdifferentiable functions in Example 4.3.
@@ 5 Subgradients in Alexandrov space
Hadamard spaces have nonpositive curvature. Loosening this restriction to allow curvature that is bounded above but possibly positive gives rise to a more general class: Alexandrov spaces. In such spaces, a notion of subgradient for convex functions appeared recently in [28]. In this section, we compare that idea with our current development. For simplicity, we restrict attention to our usual setting of a Hadamard space $X$. We briefly review some necessary tangent space constructions: for details, see [13]. For any point $x\in X$, we denote by $\Theta_{x}X$ the set of all nonconstant geodesics issuing at $x$.
@@
The Alexandrov angle induces a metric on the set $\Sigma_{x}X$ of equivalence classes of geodesics in $\Theta_{x}X$, where two geodesics $\gamma,\eta\in\Theta_{x}X$ are considered equivalent if $\angle(\gamma,\eta)=0$. For a geodesic $\gamma\in\Theta_{x}X$, we denote by $\gamma]$ its equivalence class. The tangent space $T_{x}X$ of $X$ at $x$ is the Euclidean cone over the metric space $(\Sigma_{x}X,\angle)$ (see [13, Chapter I, Definition 5.6]). For $v=([\gamma],r),w=([\eta],s)\in T_{x}X$ the scalar product of $v$ and $w$ is given by $\langle\langle v,w\rangle\rangle=rs\cos\angle(\gamma,\eta)$. For a geodesically convex function $f\colon X\to\mathbb{R}$, the subdifferential of $f$ at $x$, denoted by $\partial f(x)$, can be characterized (see [28, Proposition 4.4]) as the elements $([\eta],s)\in T_{x}X$ satisfying the subgradient inequality
@@
where $\gamma_{y}\colon[0,d(x,y)]\to X$ denotes the geodesic from $x$ to $y$. Any element $v\in\partial f(x)$ is called a subgradient of $f$ at $x$. A related idea used in the study of gradient flows in Alexandrov spaces is the minus-subdifferential [20, Definition 3.6].
@@
In general, the idea of a Busemann subgradient is more restrictive than the Alexandrov space notion. Indeed, on Alexandrov spaces, continuous geodesically convex functions have subgradients everywhere [28, Theorem 4.11], but in the special case of Hadamard spaces, we have seen (Example 3.4) that such functions may not be Busemann subdifferentiable.
As a concrete example, consider Example 4.3. The function $f$ has no Busemann subgradient at the point $\bar{x}$, but it does have a subgradient there in the Alexandrov space sense. Consider the geodesic $\eta\colon[0,1]\to X$ given by $\eta(t)=\bar{x}+t(1,1,0)$ in the quadrant $\mathbb{R}_{+}\times\mathbb{R}_{+}\times\left\{0\right\}\subseteq X$. We show in the appendix that $([\eta],3/\sqrt{2})\in T_{\bar{x}}X$ is a subgradient for $f$ at $\bar{x}$. Roughly speaking, the fact that Busemann subgradients correspond to subgradients but not conversely is due to a local/global divide: subgradients are characterized by local geometry, whereas the Busemann subgradient property is a global notion.

### v4

@@ 1 Introduction
The algorithm then tracks the geodesic ray from $x$ associated with $b$ at speed $s$ for a judiciously chosen time interval, before repeating the process. Analogous subgradient inequalities appeared in [27, 30], and indeed, the notion of a Busemann subgradient that we introduce is related to ideas in [21, 30]. The algorithms we present generalize the classical stochastic subgradient method, when choosing components uniformly at random, and the incremental subgradient algorithm of [32]: furthermore, they enjoy the same complexity. While the $p$-mean problem serves as a familiar testing ground for our algorithms, we emphasize that Busemann subgradient ideas apply more broadly: examples include the minimum enclosing ball problem [4], and, as highlighted recently in [17] [^1] While revising this paper from the original posted manuscript [22], we learned of subsequent developments involving the Busemann subgradient idea [17]., Tyler’s M-estimator [40] and Horn’s problem on deciding if given real vectors are the spectra of Hermitian matrices summing to zero [29], [28, p. 1] [^2] A concrete formulation of Horn’s problem in terms of Busemann functions appears in [17, p. 19], an approach originating from the work of [28]..
@@ 2 Geodesic geometry and convexity
We will restrict our attention to a class of functions that interacts nicely with the geometry of the given Hadamard space in a way that goes beyond geodesic convexity. Fundamental to our development is an appropriate notion of direction in a Hadamard space; in the Euclidean setting, the role played by directions and the associated compactification of $\mathbb{R}^{n}$ has been emphasized as a pillar of modern variational analysis [37, Chapter 3].
@@
We first review some standard tools to study the geometry at infinity of a Hadamard space, a more detailed discussion of which can be found in [14, Chapter II.8]. Two rays $r,r^{\prime}$ are said to be asymptotic if there exists a positive constant $K\geq 0$ such that $d(r(t),r^{\prime}(t))\leq K$ for all $t\geq 0$. This defines an equivalence relation on rays in $X$: the set of equivalence classes is denoted by $X^{\infty}$ and called the boundary of $X$ at infinity, or simply the boundary. Note our assumption that $X$ has at least two points and the geodesic extension property guarantees the boundary is nonempty. The equivalence class of a particular ray $r$ is denoted by $r(\infty)$, and we say a ray $r$ has direction $\xi\in X^{\infty}$ if $r$ belongs to the equivalence class $\xi$. Given $x\in X$ and $\xi\in X^{\infty}$, there exists a unique ray $r$ issuing from $x$ such that $r(\infty)=\xi$ [14, II.8.2]. If we fix an arbitrary reference point $\bar{x}\in X$, we can thus identify any $\xi\in X^{\infty}$ with the unique ray $r_{\bar{x},\xi}$ issuing from $\bar{x}$ with direction $\xi$, and hence to the unique corresponding Busemann function which we may denote by either $b_{r_{\bar{x},\xi}}$ or $b_{\bar{x},\xi}$. To avoid cumbersome notation, we henceforth fix such a reference point $\bar{x}\in X$ and denote the corresponding Busemann function $b_{r_{\bar{x},\xi}}$ by simply $b_{\xi}$. We note the following fact [14, Corollary II.8.20].
@@
The space $X^{\infty}$ is naturally endowed with the so-called cone topology [14, Definition II.8.6], and notably this space is first-countable [14, Proof of Theorem II.8.13]. In particular, the cone topology is completely specified by the convergent sequences, so let us mention that $\xi_{n}\to\xi\in X^{\infty}$ if and only if $b_{\xi_{n}}\to b_{\xi}$ uniformly on bounded subsets of $X$. The latter property, and hence the cone topology, is independent of the implicit basepoint (cf. [14, Proposition II.8.8]), due to Proposition 2.1.
If $X$ is proper then $X^{\infty}$ is compact [14, Definition II.8.6]. Our interest in topologizing $X^{\infty}$ comes from optimization: it will be desirable to have a large class of compact subsets of $X^{\infty}$ so that certain continuous functions defined on $X^{\infty}$ will attain their maximum. In the literature, the set $X^{\infty}$ is often endowed with a stronger metric topology induced by an angular metric (see for example [14, Proposition II.9.7], [24]). This angular metric is convenient for the study of large-scale geometry of the space $X$ but is often too strong to be useful for our purposes. For example, the boundary at infinity of the $n$-dimensional hyperbolic space $\mathbb{H}^{n}$ is discrete in this angular metric, whereas it is homeomorphic to $\mathbb{S}^{n-1}$ in the cone topology (see Example 2.3). In particular, the only compact sets are finite when the angular metric is used. Henceforth we will only make use of the cone topology.
Convergence in $X^{\infty}$ can be understood more geometrically via pointwise convergence of rays, a fact stated in [18, p.7] without proof, and also noted in [5, p.29]. The next proposition is a precise formulation: for completeness, we include a proof in the appendix. The proof and result resemble [14, Proposition II.8.19], but that result concerns unbounded sequences in $X$ that converge to points in $X^{\infty}$ whereas we emphasize sequences converging in $X^{\infty}$ itself.
@@ 3 Busemann subgradients and envelopes
is the Busemann function for the ray $r$. The argument on the right-hand side is nonincreasing in $t$, by [14, Lemma II.8.18(1)], and equals d(y,a)d(x,a)d(y,a)-d(x,a) when $t=d(x,a)$, so the result follows. ∎
@@
We consider $X=\mathbb{H}^{2}$, The subset $F=\left\{(x,y)\in\mathbb{H}^{2}\mid y=0\right\}$ is closed and geodesically convex (it is a geodesic line in $\mathbb{H}^{2}$), so the function $f(p)=\operatorname{dist}(p,F)$ is geodesically convex ([14, II.2 Corollary 2.5]). However, $f$ is not Busemann subdifferentiable on any neighborhood of $F$. To see this, let $U$ be a neighborhood of $F$, consider any $(x_{0},y_{0})\in U$ with $y_{0}\neq 0$, and suppose there exists a Busemann subgradient $\xi,s]$ at $(x_{0},y_{0})$, with $s>0$ because $(x_{0},y_{0})$ does not minimize $f$. We can identify $\xi\in X^{\infty}$ with a point $\xi=(\xi_{1},\xi_{2})\in\mathbb{S}^{1}$, and we will use this to write the Busemann function $b_{\xi}$ explicitly. By definition $f-\left\langle\cdot,[\xi,s]\right\rangle$ is minimized at $(x_{0},y_{0})$, i.e.
@@
Choosing our reference point to be $(\bar{x},\bar{y})=(0,0)$, [14, II.8.24(2)] tells us the Busemann function $b_{\xi}$ has the following form:
@@
(Continuous-compact-representable functions) Endowing the boundary $X^{\infty}$ with the cone topology opens up the following natural construction. Suppose $K\subseteq CX^{\infty}$ is compact and $g\colon CX^{\infty}\to(-\infty,+\infty]$ is continuous on $K$ and $\infty$ outside of $K$. Because the pairing is also continuous (Proposition 2.3(1)), the function $g^{\circ}$ is a Busemann envelope, and hence Busemann subdifferentiable by Proposition 3.4. This construction generates a rich class of examples. In particular, in Euclidean space we recover the class of “almost sublinear” functions: those functions $f$ such that $\sup_{x\in\mathbb{R}^{n}}|f(x)-h(x)|<\infty$ for some sublinear function $h$ [12, Proposition 4.5]. This class covers all convex optimization problems with compact feasible regions, by extending the objectives to have this growth property outside the feasible region. More generally, consider the case when the space $X$ is proper, and hence the boundary $X^{\infty}$ is compact. Then compact subsets of the boundary cone $CX^{\infty}$ are abundant, because if $F\subseteq X^{\infty}$ is closed and $I\subseteq\mathbb{R}_{+}$ is compact, then the image of $F\times I$ under the quotient map is compact.
@@
which holds with equality if and only if $\xi,s]$ is a Busemann subgradient for $f$ at $x$, analogously to the classical case. The only difference between the definitions of $f^{*}$ and $f^{\bullet}$ is the sign of the pairing term. For comparison, [9, Definition 1] concerns a notion of Fenchel conjugate on Hadamard manifolds with the sign on the pairing term matching inequality (14).
@@ 4 The geometry of Busemann subdifferentiability
Note $s_{k}>0$ because $x_{k}$ does not minimize $f$, hence $b_{x_{k},\xi_{k}}(y)\leq 0$ for all $y\in F$. Properness of $X$ implies that $X^{\infty}$ is compact, allowing us to extract a convergent subsequence $\xi_{k}\to\xi$. The map $(x,\xi)\mapsto b_{x,\xi}(y)$ is continuous [5, II.1] so passing to the limit implies $b_{x,\xi}(y)\leq 0$ for $y\in F$. Thus $F$ admits a supporting horoball at each boundary point as desired. ∎
@@
Just as in Euclidean space, continuous convex functions on the space $X$ must be locally Lipschitz, by [30, Lemma 4.3]. The following result relates Busemann subgradients to the local Lipschitz constant: for an analogous Euclidean result, see [8, Theorem 3.61].
@@
(ii) Suppose $f$ is $L$-Lipschitz on $C$ and take any Busemann subgradient $\xi,s]$ at $x\in C$. If $s=0$ there is nothing to prove, so we assume $s>0$. Let $r$ be the ray issuing from $x$ with direction $\xi$. Since $X$ has the geodesic extension property $r$ can be extended to a geodesic line $\tilde{r}\colon\mathbb{R}\to X$ by [14, Lemma II.5.8(2)]. The geodesic line $\tilde{r}$ is continuous with $\tilde{r}(0)=x\in C$, and $C$ is open so there exists $\varepsilon>0$ such that $\tilde{r}(-\varepsilon)\in C$. By (6) we have
@@
This space arises as a subspace of the tree space $\mathcal{T}_{4}$ (see Section 8), and is illustrated in [11, Figure 12] as well as Figure 4.1. We will revisit this space when discussing the Busemann subdifferentiability of a sum of Busemann subdifferentiable functions in Example 4.3.
@@ 5 Subgradients in Alexandrov space
Despite its generality, our development in this work is limited to the nonpositively-curved setting of Hadamard spaces. Loosening this significant restriction to allow curvature that is bounded above but possibly positive gives rise to a more general class: Alexandrov spaces. Convexity of distance functions then only holds on domains of small diameter. (Euclidean spheres, for example, require restricting to small spherical caps.) Nonetheless, on such domains, $p$-mean problems (for example) can be solved via Riemannian gradient descent, on manifolds [1], and more generally via proximal methods [33]. However, beyond Hadamard spaces, Busemann subgradients seem unhelpful: geodesic rays do not exist in spheres, for example. Whether some other notion of subgradient might support optimization algorithms on Alexandrov spaces with a satisfactory complexity analysis remains an open question. As background, we here compare Busemann subgradients with a recent notion of subgradient for convex functions on Alexandrov spaces [30]. For simplicity, we restrict attention to our usual setting of a Hadamard space $X$. We briefly review some necessary tangent space constructions: for details, see [14]. For any point $x\in X$, we denote by $\Theta_{x}X$ the set of all nonconstant geodesics issuing at $x$.
@@
The Alexandrov angle induces a metric on the set $\Sigma_{x}X$ of equivalence classes of geodesics in $\Theta_{x}X$, where two geodesics $\gamma,\eta\in\Theta_{x}X$ are considered equivalent if $\angle(\gamma,\eta)=0$. For a geodesic $\gamma\in\Theta_{x}X$, we denote by $\gamma]$ its equivalence class. The tangent space $T_{x}X$ of $X$ at $x$ is the Euclidean cone over the metric space $(\Sigma_{x}X,\angle)$ (see [14, Chapter I, Definition 5.6]). For $v=([\gamma],r),w=([\eta],s)\in T_{x}X$ the scalar product of $v$ and $w$ is given by $\langle\langle v,w\rangle\rangle=rs\cos\angle(\gamma,\eta)$. For a geodesically convex function $f\colon X\to\mathbb{R}$, the subdifferential of $f$ at $x$, denoted by $\partial f(x)$, can be characterized (see [30, Proposition 4.4]) as the elements $([\eta],s)\in T_{x}X$ satisfying the subgradient inequality
@@
where $\gamma_{y}\colon[0,d(x,y)]\to X$ denotes the geodesic from $x$ to $y$. Any element $v\in\partial f(x)$ is called a subgradient of $f$ at $x$. A related idea used in the study of gradient flows in Alexandrov spaces is the minus-subdifferential [21, Definition 3.6].
@@
In general, the idea of a Busemann subgradient is more restrictive than the Alexandrov space notion. Indeed, on Alexandrov spaces, continuous geodesically convex functions have subgradients everywhere [30, Theorem 4.11], but in the special case of Hadamard spaces, we have seen (Example 3.4) that such functions may not be Busemann subdifferentiable.
As a concrete example, consider Example 4.3. The function $f$ has no Busemann subgradient at the point $\bar{x}$, but it does have a subgradient there in the Alexandrov space sense. Consider the geodesic $\eta\colon[0,1]\to X$ given by $\eta(t)=\bar{x}+(t/\sqrt{2})(1,1,0)$ in the quadrant $\mathbb{R}_{+}\times\mathbb{R}_{+}\times\left\{0\right\}\subseteq X$. We show in the appendix that $([\eta],3/\sqrt{2})\in T_{\bar{x}}X$ is a subgradient for $f$ at $\bar{x}$. Roughly speaking, the fact that Busemann subgradients correspond to subgradients but not conversely is due to a local/global divide: subgradients are characterized by local geometry, whereas the Busemann subgradient property is a global notion.